Circular No. 8567

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2005cw IN IC 1439 Y.-T. Chen, Graduate Institute of Astronomy, National Central Uni-

versity; together with T.-W. Chen and Z.-Y. Lin, report the discovery of an apparent supernova (mag ∼ 18.5) on unfiltered CCD images taken on July 11.853 UT with the 1-m Lulin Observatory telescope in the course of the Taiwan Supernova Survey; the new object was confirmed at mag ≈ 18.1 on an image taken on July 12.651. SN 2005cw is located at α = 22^{h}16^{m}40^{s}.14, δ = −21^{o}29′08′′.9 (equinox 2000.0), which is 5′′.6 west and 20′′.5 south of the center of IC 1439. Nothing was visible at this location on an image taken on 2004 Dec. 14.510 (limiting magnitude ∼ 18.6).

AS 325 S. Otero, Buenos Aires, Argentina, reports that AS 325, an eclipsing

binary (period 513 days), has entered a rapid decline phase and is currently near mid-eclipse; the current eclipse started on June 5, and eclipses last ∼ 86 days. AS 325 is presently at V = 12.0, or 2 mag fainter then prior to eclipse, and over 1 mag fainter than during previous eclipses (IBVS

5608; http://www.astrouw.edu.pl/cgi-asas/asas lc/185004-2624.3).

S. Howell, WIYN Observatory, reports that spectra obtained during June 28–July 2 reveal that the H emission is weaker by 50 percent, while the Fe II emission remains constant, and the previously strong Ca II H and K absorption is nearly absent (Cool et al. 2005, PASP 117, 462). Spectra from July 1 show a slightly increasing blue continuum mainly consisting of emission lines and obvious TiO absorption in the red. Mid-eclipse is expected for July 18, and multicolor observations and additional spectra are needed to confirm that the cool component is the one that is being eclipsed. The Tycho Reference Catalogue (via SIMBAD) lists the position of AS 325 as α = 18^{h}50^{m}03^{s}.57, δ = −26^{o}24′15′′.4 (equinox 2000.0).

TIME ADJUSTMENT ON 2005 DECEMBER 31 Bulletin C30 of the International Earth Rotation and Reference Sys-

tems Service announces that a positive leap second will be introduced such that the sequence of UTC second markers will be: 2005 Dec. 31^{d}23^{h}59^{m}59^{s}, 31^{d}23^{h}59^{m}60^{s}, 2006 Jan. 1^{d}00^{h}00^{m}00^{s}. Beginning 2006 Jan. 1, the difference UTC–TAI = −33 s (UTC–TAI has been −32 s since 1999 Jan. 1).

2005 July 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green